OpenStudy (anonymous):

the distance EA for the points E(2, 3) and A(4, 2)

OpenStudy (anonymous):

the distance is \[\sqrt{5}\] by using the formula \[\sqrt{(4-2)^2+(1-2)^2}\]

OpenStudy (anonymous):

no its \[\sqrt{(4-2)^2+(2-3)^2}\]

OpenStudy (anonymous):

which is \[\sqrt{5}\]